# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

### 05 Feb 2024 v1

@@ 2.3. Leveraging AI to support writing
Recent works explored the potential of using AI to support people’s workflow on writing tasks (Ippolito et al., 2022; Chung et al., 2022; Clark et al., 2018; Gero and Chilton, 2019b, a; Rahman et al., 2020; Lee et al., 2022; Gero et al., 2021; Ippolito et al., 2022; Mirowski et al., 2022). For example, TaleBrush allows users to create a story with AI through sketching to aid the planning of writing. Then it enables writers to generate diverse storylines and interactively refine them (Chung et al., 2022). Another system Wordcraft explores how to support users collaborating with generative language models to co-write a story (Yuan et al., 2022). Spark used a language model to generate prompts related to a scientific concept to facilitate scientific writing (Gero et al., 2021).
@@
Research showed that writing is a complex, iterative process (Flower and Hayes, 1981). The Hayes model describes the cognitive processes an individual writer engages in during the process of writing (Flower and Hayes, 1981; Hayes, 2012; Gero et al., 2022). In the writing process, there are three major components: the planning component, the translating component, and the reviewing component, as shown in Figure 5. To facilitate the cognitive process of writing, several systems are developed to support different stages (Bhat et al., 2023; Zhang et al., 2023). For example, VISAR is an AI-enabled writing assistant to helps writers brainstorm and revise hierarchical goals and organize argument structures in the planning stage.
@@
To facilitate the iterative planning and revising process in writing, intelligent systems further use the chain of thoughts (COT) prompting method to break down the large problem into step-by-step prompts (Wei et al., 2022). Specifically, Re3 framework and DOC framework used the COT approach to decompose the writing tasks where they first generate an outline and then automatically turn the outline into the story generation. The evaluation demonstrated that this decomposition approach can improve the coherence of long story generation (Yang et al., 2022b, a). However, this approach could be highly controllable where humans can control the story generation by modifying the outlines. Control and agency are extremely important in the conference peer review writing where human reviewers should play the role of driving the writing process. The development of LLM brings opportunities along with concerns on what LLMs shouldn’t do in the academic writing area (Fok and Weld, 2023). The prior work that aligns most closely with the concept of applying AI techniques to reviewing academic papers that a machine model that automatically generates feedback using LLMs (Liang et al., 2023). Results showed that LLM feedback could benefit researchers in earlier stages of manuscript preparation while researchers struggle with an in-depth critique of study methods. Instead of using an automatic method, humans should be in the loop to drive and control the writing process (Fok and Weld, 2023). Drawing on these insights, we designed the ReviewFlow system not to automate any parts of the process, but rather to scaffold key considerations and to give novices agency over how to apply machine-generated language suggestions.
@@ 3. Formative Studies
Before we developed our system to support academic peer reviewing, we conducted a series of formative studies. First, we interviewed ten novice reviewers to understand how they approached this task for the first time and their perceived challenges(represented as C). Second, we observed ten experienced reviewers as they conducted a peer review on a couple of selected papers to understand common practices and workflows. Last, we propose design goals((represented as DG)) on supporting novice reviewers.
@@ 3.1.1. Novice Interview Study
We conducted 30-minute semi-structured interviews with 10 novice reviewers with relatively little experience with writing academic peer reviews (ranging from only reviewing once before and having less than 2 years of experience). Participants were recruited through mailing lists and social media posts. The participants (4 female and 6 male, average age of 25.5 years) came from diverse research fields including Human-Computer Interaction, Artificial Intelligence, Cognitive Science, and Computer Science. Interviews were conducted remotely by the lead author and lasted around 30 minutes.
In the interview, we first asked open-ended questions about their perceived obstacles and challenges of conducting academic peer review. We then provided scenarios of potential features for peer review scaffolding to elicit reactions from novice reviewers. These scenarios described common potential situations faced by first-time reviewers and used these to elicit participants’ needs in a real-world situation (Gaver et al., 1999). For example, “Mary is a first-year graduate student doing Human-Computer Interaction (HCI) research. While she has submitted a couple of papers before she received reviews, this will be her first time as a reviewer. She struggled to make the review constructive for the authors.”. A complete list of scenarios and design probes is listed in the Appendix. Based on these answers, three researchers on our team iterated multiple times to come up with design probes targeting these challenges and visualize them as design probes in Figma.
In the end, we provided a series of design probes for the novices to capture their reactions. Drawing insights from previous literature around the intelligent support on reading and writing, the research team took multiple rounds of iteration and group discussions to develop the six design probes (Zhang et al., 2023; Lo et al., 2023). The design probes included: (1) scaffold annotation with community curated tags and filters, (2) contextual reflection cues (expert-authored generic cues versus AI-generated specific reflection cues on each section or highlighted sentences), (3) extractive summarization and generated explanations to facilitate paper reading, (4) in-situ citation recommendation where the system provides a summary from the cited paper and recommend potential missing citations, (5) review draft generation with structures, (6) mapping back the source of review draft from the paper content and visualize the location of the source to help reviewers revise their draft. Interviews were recorded with participants’ permission and were transcribed into text. Three researchers from the team went through the transcripts and coded them for themes using thematic analysis (Braun and Clarke, 2006). Through multiple iterations along with periodic discussions with the rest of the research team, the coding led to the major themes of challenges below.
@@ 3.1.2. Experts Observational Study
We also chose to recruit experienced peer reviewers to learn about their best practices through a contextual inquiry (Karen and Sandra, 2017). We recruited 10 experts with at least 5 years of experience in writing academic peer reviews. Participants were recruited through in-person invitations, email lists, and social media posts. Participants (4 female and 6 male, average age of 30.1 years) came from different fields of computer science, including Human-Computer Interaction, Artificial Intelligence, programming languages, learning science, computer security, and accessibility. Observational studies were conducted remotely by the lead author and lasted around 90 minutes.
During the observational study, we first spent 20 minutes asking about their experiences and challenges with peer reviewing. Then, the team asked the participant to write a peer review in a Google document for one of 5 different short papers (less than 4 pages) in 60 minutes. To find suitable short papers, the team first collected the participants’ research interest descriptions from their websites and used these keywords to search on Semantic Scholar (sem, 2015). We filtered out the papers longer than 5 pages, ranked them by relevance, and selected the top N as the participant’s options for the peer review exercise. Last, we provided the same design probes to the experts to elicit what scaffolding strategies could benefit the novices the most. The interviews were recorded and then transcribed using a machine transcription service. The research team took the same analysis procedure as the previous novice study.
@@ 3.2. Findings
3.2.1. Novice reviewers felt they lacked guidance on evaluating the paper and writing structured and constructive reviews
The novice participants reported that their prior attempts at reviewing papers were cognitively demanding and took an average of 6.4 hours. When asked about challenges, 6/10 mentioned that they struggle to fully understand the background knowledge and existing work and did not feel confident about assessing the paper’s novelty [C2]. 4/10 mentioned that they lacks opportune considerations for assessment [C1]. Specifically, one participant mentioned that they would love to have “co-reviewers who have already read the paper and known the specifics to guide me through the evaluation process”(N5). In addition, 4/10 highlighted that they need more guidance on how to structure issues during the writing process[C3]. 3/10 explicitly mentioned that they want to receive some feedback from experts, other reviewers, or authors, to make sure the review is in the right tone and meet the expectation [C4]. Specifically, one participant mentioned “I am not sure whether I covered all the necessary points or whether authors will perceive the reviews as useful”(N7). Another participant is worried about being “impolite or harsh (N9)”.
We presented novices the design probes (see Appdendix) one by one and asked them to rate the degree to which they resonate with each scenario (1=Not resonate at all to 5= Strongly resonate with the situation). The top two situations that resonate with novice reviewers the most were: “the novice reviewer struggles to write a high-quality review”(3.8/5) and “the novice reviewer spends a lot of time reading a paper but doesn’t know how to evaluate the paper”(3.8/5).
@@ 3.2.2. Experienced reviewers adopt a workflow of sense-making, annotating, and synthesizing notes
After finishing reading the paper, instead of directly editing the review draft, experts created high-level headers or topics, such as “lack clarity on study design”, that summarized the paper’s weaknesses and strengths as well as prioritized the concerns (7/10). For instance, one expert reflected on the review process as “I will make a lot of annotations and then I will just review them section by section. At the same time, thinking about what is the biggest concern in terms of the overall research novelty”(E6). We observed that 6 out of 7 participants wrote the review in the following structure: summary of the paper, strengths or contribution of the paper, two to three weaknesses of the paper, and end the review with decision justification and general recommendations. Some experts will then list bullet points under each topic together with questions that they would like to ask the author. Last, experts compiled these comments and bullet points into a complete draft and revised them two to three times to offer suggestions and make it more constructive. Figure 2 represents the expert’s review workflow together with their practices.
@@ 3.2.3. Experienced reviewers stressed the importance of specific and contextual guidance for scaffolding
Among these features, 5 experts reflected that the contextual cues can be helpful, especially for novice reviewers. One participant preferred the AI-generated cues on the selected paper content and explained that “I like to have some capability of freedom, but I think this will be super useful for novice reviewers who don’t know how to review. But for people who have reviewed for so many years, those guidances for conferences are pretty much the same”(E6). We provided experts with two sets of tags to select. The first set of tags is designed based on the review structure that includes “summary of the paper”, “strength”, “weakness” and “others”. The second set of tags is designed using community review criteria that include “relevance”, “novelty”, “validity”, “clarity”. Most experts preferred the first set of tags (7/10). 2 participants mentioned that they were concerned the experts’ authored cues might be too similar to the existing guidelines. Hence, they suggested the contextual question can provide better guidance for novices.
9 out of 10 experts mentioned their preference for the in-situ citation support to provide summary and recommendations. They reflected that this in-situ knowledge support can “raise awareness on unknown work”(E5). 9 out of 10 participants expressed their concerns about the summarization feature where they don’t trust the AI’s ability to identify important information. Instead, they think reviewers should have control over the reading process. Specifically, one expert mentioned “I think we are also reviewing the style of writing, or how something is communicated and logically connected between each paragraph or each sentence. So I think there is value with actually reading everything to get the message behind the paragraphs”(E3). Participants shared their opinions on using AI in the review process and all of them agreed that LLM should not generate the review on the fly while human experts should drive the process since the limitation of LLM can bias human experts and bring over-reliance on the use of AI.
Based on the interview on the design probes, we summarize the following design considerations mentioned by novices and experts. The support should be lightweight and don’t distract the current review flow. While involving AI in the review process, AI should not go too far to biases or lead the thinking process. Human reviewers should still preserve agency in reading, writing, and decision-making. Faced with the limitations of current LLM, intelligent systems should try to avoid hallucinations and provide enough opportunities for fact-checking.
@@ 3.3. Formative Study Discussion
DG1: Offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper Novices highlighted their need for guidance in critically evaluating papers from various perspectives, expressing uncertainty about the key points to focus on [C1]. In educational settings, instructors typically serve as scaffolds when introducing new concepts or knowledge to novice learners (Collins, 2006). Prior studies in scientific paper reading developed methods to  (August et al., 2022; Rachatasumrit et al., 2022). For example, Paper Plain provided a collection of key questions that guide readers to answering passages and plain language summaries of those passages (August et al., 2022). Similarly, in the context of our study, the objective is to provide guidance comparable to that of expert peer reviewers. The emphasis is on delivering locally relevant questions that help users think critically about each section of the paper.
@@
DG2: Enable in-situ knowledge support for assessing novelty compared to prior work Novices are struggled with insufficient background knowledge for evaluating papers [C2]. Providing knowledge support in situ can externalize the user’s working memory, aid in sense-making and facilitate a swift reviewing and resumption of task contexts (Kuznetsov et al., 2022). Prior studies in scientific literature review highlight the importance of in-situ knowledge support (Kang et al., 2022; Chang et al., 2023). To help novice reviewers who lack background knowledge to evaluate the novelty while reading the paper, our goal is to enable in-situ knowledge support by extracting the abstract of the cited paper.
@@
DG3: Model the expert workflow of reading, evaluating and synthesizing practiced by experienced reviewers Novices expressed their need for more specific guidance from experts in terms of their process and structure [C3]. Prior work showed that surfaced expert practices can better structure and scaffold the process for novices (Kim et al., 2015). Motify illustrates that storytelling patterns extracted from expert stories can be used to effectively scaffold novices to create video stories (Kim et al., 2015). In the context of writing introductory help requests, providing high-quality examples and expert-informed templates can increase learning and writing quality (Hui et al., 2018). Our approach involves the modeling of the expert workflow encompassing reading, evaluating, synthesizing, and revising. This model is then employed to structure the peer review process for novice reviewers.
@@
DG4: Guide the reviewing process such that the work aligns with community standards Novices reported that they lacked effective models for tone and expectation [C4]. More guidance on the expectation of the review can help reviewers to reflect on their tone and structure. In addition, experts raised valid concerns about the potential for intelligent tools to introduce biases or influence the thinking process. Existing studies have highlighted apprehensions regarding AI exhibiting biases in human-AI collaboration tasks (Kawakami et al., 2022; Poursabzi-Sangdeh et al., 2021). Notably, research has shown that providing information, including explanations, generated by AI has the potential to mislead users in decision-making (Lakkaraju and Bastani, 2020; Goyal et al., 2023; Buçinca et al., 2021). Therefore, a key design objective is to steer clear of biasing the decision-making process and focus solely on encouraging justifications. This approach aims to assist novice reviewers in producing high-quality and original reviews without compromising the integrity of the evaluation process.

### 26 Feb 2024 v2

@@ 2.3. Leveraging AI scaffolding to support writing
Advances in AI and LLMs provide opportunities to provide efficient and context-specific scaffolding in the learning process (Ippolito et al., 2022; Chung et al., 2022; Clark et al., 2018; Gero and Chilton, 2019b, a; Rahman et al., 2020; Lee et al., 2022; Gero et al., 2021; Ippolito et al., 2022; Mirowski et al., 2022). For example, TaleBrush allowed users to create a story with AI through sketching to aid the planning of writing. Then it enabled writers to generate diverse storylines and interactively refine them (Chung et al., 2022). Another system Wordcraft explored how to support users collaborating with generative language models to co-write a story (Yuan et al., 2022). Spark used a language model to generate prompts related to a scientific concept to facilitate scientific writing (Gero et al., 2021). However, as far as we know, none of the studies that used scaffolding strategies focused on the context of conference peer review writing. Our research explores how we might guide a complex, multi-faceted workflow like peer reviewing and the potential role of AI in creating contextually adaptive scaffolds.
@@
Writing is a complex, iterative process (Flower and Hayes, 1981). The Hayes model describes the cognitive processes an individual writer engages in during the process of writing (Flower and Hayes, 1981; Hayes, 2012; Gero et al., 2022). In the cognitive process of writing, there are three major components: planning, translating, and reviewing, as shown in Figure 5. Several systems are developed to facilitate different stages of writing (Bhat et al., 2023; Zhang et al., 2023). For example, VISAR is an AI-enabled writing assistant that helps writers brainstorm and revise hierarchical goals and organize argument structures in the planning stage. To facilitate the iterative planning and revising process in writing, intelligent systems further use the chain of thoughts (COT) prompting method to break down the large problem into step-by-step prompts (Wei et al., 2022). Specifically, the Re3 framework and DOC framework used the COT approach to decompose the writing tasks where they first generate an outline and then automatically turn the outline into the story generation (Yang et al., 2022b, a). Their evaluation demonstrated that the decomposition approach can improve the coherence of long story generation and is highly controllable where humans can control the story generation by modifying the outlines.
@@
Control and agency are extremely important in the conference peer review writing process, as human reviewers should play the role of driving the writing process. While the development of LLMs can provide scaffolding opportunities for this writing process, it is important to address the concerns and limitations of LLMs in this context (Fok and Weld, 2023). The prior work that aligns most closely with the concept of applying AI techniques to reviewing academic papers that a machine model that automatically generates feedback using LLMs. Results showed that LLM feedback could benefit researchers in earlier stages of manuscript preparation while researchers struggle with an in-depth critique of study methods (Liang et al., 2023). Instead of using an automatic method, humans should be in the loop to drive and control the writing process (Fok and Weld, 2023). Drawing on these insights, we designed the ReviewFlow system not to automate any parts of the process, but rather to scaffold key considerations and to give novices agency over how to apply machine-generated language suggestions.
@@ 3. Formative Studies
Before we developed our system to support academic peer reviewing, we conducted two formative studies. First, we interviewed ten novice reviewers to understand how they approached this task for the first time and their perceived challenges (represented as C). Second, we conducted observational studies with ten experienced reviewers where we invited them to write a peer review on a selected paper to capture their common practices and workflows. Last, based on the findings, we proposed design goals (represented as DG) for supporting novice reviewers.
@@ 3.1.1. Novice Interview Study
We conducted semi-structured interviews with 10 novice reviewers (N1-N10) with relatively little experience with writing academic peer reviews (ranging from only reviewing once before and having less than 2 years of experience). Participants were recruited through mailing lists and social media posts. The participants (4 female and 6 male, average age of 25.5 years) came from diverse research fields including Human-Computer Interaction, AI, Cognitive Science, and Computer Security. Interviews were conducted remotely by the lead author and lasted around 30 minutes.
In the interview, we first asked open-ended questions about their perceived obstacles and challenges of conducting academic peer review. We then provided scenarios of potential features for peer review scaffolding to elicit reactions from novice reviewers. These scenarios described common potential situations faced by first-time reviewers and were designed to prompt the participants to share their needs in a real-world situation (Gaver et al., 1999). For example, “Mary is a first-year graduate student doing Human-Computer Interaction (HCI) research. While she has submitted a couple of papers before she received reviews, this will be her first time as a reviewer. She struggled to make the review constructive for the authors”. Interviews were recorded with participants’ permission and were transcribed. Two researchers from the team went through the transcripts and coded themes using thematic analysis (Braun and Clarke, 2006). Through multiple iterations along with periodic discussions, the coding led to the major themes of challenges below.
@@ 3.1.2. Experts Observational Study
We conducted observational studies with 10 experts reviewers(E1-E10) with at least 5 years of experience in writing academic peer reviews, to learn about their best practices (Karen and Sandra, 2017). Participants were recruited through in-person invitations, email lists, and social media posts. Participants (4 female and 6 male, average age of 30.1 years) came from different fields of computer science, including HCI, AI, programming languages, learning science, computer security, and accessibility. Observational studies were conducted remotely by the lead author and lasted around 90 minutes.
During the observational study, we first spent 20 minutes asking about their experiences and challenges with peer reviewing. Then, the team asked the participant to write a peer review in a Google document for one of 5 different short papers (less than 4 pages) in 60 minutes. To find suitable short papers, the team first collected the participants’ research interest descriptions from their websites and used these keywords to search on Semantic Scholar (sem, 2015). We filtered out the papers longer than 5 pages, ranked them by relevance, and selected the top 5 as the participant’s options for the study.
In the end, we provided a series of design probes for the novices to capture their reactions. Drawing insights from previous literature around the intelligent support on reading and writing, three researchers on our team took multiple rounds of iteration and group discussions to develop the six design probes (Zhang et al., 2023; Lo et al., 2023). These design problems are visualized in Figma. The design probes included: (1) scaffold annotation with community curated tags and filters, (2) reflection questions (expert-authored generic questions versus AI-generated specific reflection questions on each section or highlighted sentences), (3) extractive summarization and generated explanations to facilitate paper reading, (4) in-situ citation recommendation where the system provides a summary from the cited paper and recommend potential missing citations, (5) review draft generation, (6) mapping back the source of review draft from the paper content and visualize the location of the source to help reviewers revise their draft. The interviews were recorded and then transcribed using a machine transcription service. The research team took the same analysis procedure as the novice study.
@@ 3.2. Findings
3.2.1. Novice reviewers felt they lacked guidance in evaluating the paper and writing structured and constructive reviews
The novice participants reported that their prior attempts at reviewing papers were cognitively demanding and took an average of 6.4 hours. When asked about challenges, 6/10 mentioned that they struggled to fully understand the background knowledge and existing work and did not feel confident about assessing the paper’s novelty (C2). 4/10 mentioned that they lack opportune considerations for assessment (C1). Specifically, one participant mentioned that they would love to have “co-reviewers who have already read the paper and known the specifics to guide me through the evaluation process”(N5). In addition, 4/10 highlighted that they need more guidance on how to structure issues during the writing process(C3). 3/10 explicitly mentioned that they want to receive some feedback from experts, other reviewers, or authors, to make sure the review is in the right tone and meets the expectation (C4). Specifically, one participant mentioned “I am not sure whether I covered all the necessary points or whether authors will perceive the reviews as useful”[N7]. Another participant is worried about being “impolite or harsh”[N9].
We presented novices with the design scenarios one by one and asked them to rate the degree to which they resonate with each scenario (1 = Not resonate at all to 5 = Strongly resonate with the situation). The top two situations that resonated with novices the most were: “the novice reviewer struggles to write a high-quality review” (3.8/5) and “the novice reviewer spends a lot of time reading a paper but doesn’t know how to evaluate the paper” (3.8/5).
@@ 3.2.2. Experienced reviewers adopt a workflow of sense-making, annotating, and synthesizing notes
After finishing reading the paper, instead of directly editing the review draft, experts created high-level headers or topics, such as “lack clarity on study design”, that summarized the paper’s weaknesses and strengths as well as prioritized the concerns (7/10). For instance, one expert reflected on the review process as “I will make a lot of annotations and then I will just review them section by section. At the same time, thinking about what is the biggest concern in terms of the overall research novelty”(E6). We observed that 7 participants wrote the review in the following structure: summary of the paper, strengths or contribution of the paper, two to three weaknesses of the paper, and end the review with decision justification and general recommendations. Some experts will then list bullet points under each topic together with questions that they would like to ask the author. Last, experts compiled these comments and bullet points into a complete draft and revised them two to three times to offer suggestions and make it more constructive. Figure 2 represents the expert’s review workflow together with their practices.
@@ 3.2.3. Experienced reviewers stressed the importance of specific and contextual guidance for scaffolding
Among these features, 5 experts reflected that the contextual cues can be helpful, especially for novice reviewers. One participant preferred the AI-generated cues on the selected paper content and explained that “I like to have some capability of freedom, but I think this will be super useful for novice reviewers who don’t know how to review. But for people who have reviewed for so many years, those guidances for conferences are pretty much the same” [E6]. We provided experts with two sets of tags to select. The first set of tags is designed based on the review structure that includes “summary of the paper”, “strength”, “weakness” and “others”. The second set of tags is designed using community review criteria that include “relevance”, “novelty”, “validity”, “clarity”. Most experts preferred the first set of tags (7/10). 2 participants mentioned that they were concerned the experts’ authored cues might be too similar to the existing guidelines. Hence, they suggested the contextual question can provide better guidance for novices.
9 out of 10 experts mentioned their preference for the in-situ citation support to provide summary and recommendations. They reflected that this in-situ knowledge support can “raise awareness on unknown work” [E5]. 9 out of 10 participants expressed their concerns about the summarization feature, as they don’t trust the AI’s ability to identify important information. Instead, they think reviewers should have control over the reading process. Specifically, one expert mentioned “I think we are also reviewing the style of writing, or how something is communicated and logically connected between each paragraph or each sentence. So I think there is value with actually reading everything to get the message behind the paragraphs” [E3]. Participants shared their opinions on using AI in the review process and all of them agreed that LLMs should not generate the review on the fly, and instead, human experts should drive the process since the limitation of LLMs can bias human experts and lead to over-reliance on the use of AI.
Based on the interview on the design probes, we summarize the following design considerations mentioned by novices and experts. The support should be lightweight and not distract from the current review flow. While involving AI in the review process, AI should not go too far to bias or lead the thinking process. Human reviewers should still preserve agency in reading, writing, and decision-making. Faced with the limitations of current LLMs, intelligent systems should try to avoid hallucinations and provide enough opportunities for fact-checking.
@@ 3.3. Formative Study Discussion
DG1: Offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper Novices highlighted their need for guidance in critically evaluating papers from various perspectives, expressing uncertainty about the key points to focus on (C1). In educational settings, instructors typically serve as scaffolds when introducing new concepts or knowledge to novice learners (Collins, 2006). Prior studies in scientific paper reading developed methods to offer contextual hints or guided questions that encouraged reflection and critical thinking (August et al., 2022; Rachatasumrit et al., 2022). For example, Paper Plain provided a collection of key questions that guide readers to answering passages and plain language summaries of those passages (August et al., 2022). Similarly, in the context of our study, the objective is to provide guidance comparable to that of expert peer reviewers. The emphasis is on delivering locally relevant questions that help users think critically about each section of the paper.
@@
DG2: Enable in-situ knowledge support for assessing novelty compared to prior work Novices struggled with insufficient background knowledge for evaluating papers (C2). Providing knowledge support in situ can externalize the user’s working memory, aid in sense-making, and facilitate a swift reviewing and resumption of task contexts (Kuznetsov et al., 2022). Prior studies in scientific literature review highlight the importance of in-situ knowledge support (Kang et al., 2022; Chang et al., 2023). To help novice reviewers who lack background knowledge to evaluate the novelty while reading the paper, our goal is to enable in-situ knowledge support by extracting the abstract of the cited paper.
@@
DG3: Model the expert workflow of reading, evaluating, and synthesizing practiced by experienced reviewers Novices expressed their need for more specific guidance from experts in terms of their process and structure (C3). Prior work showed that surfaced expert practices can better structure and scaffold the process for novices (Kim et al., 2015). Motify illustrates that storytelling patterns extracted from expert stories can be used to effectively scaffold novices to create video stories (Kim et al., 2015). In the context of writing introductory help requests, providing high-quality examples and expert-informed templates can increase learning and writing quality (Hui et al., 2018). Our approach involves the modeling of the expert workflow encompassing reading, evaluating, synthesizing, and revising. This model is then employed to structure the peer review process for novice reviewers.
@@
DG4: Guide the reviewing process such that the work aligns with community standards Novices reported that they lacked effective models for tone and expectation (C4). More guidance on the expectation of the review can help reviewers to reflect on their tone and structure. In addition, experts raised valid concerns about the potential for intelligent tools to introduce biases or influence the thinking process. Existing studies have highlighted apprehensions regarding AI exhibiting biases in human-AI collaboration tasks (Kawakami et al., 2022; Poursabzi-Sangdeh et al., 2021). Notably, research has shown that providing information, including explanations, generated by AI has the potential to mislead users in decision-making (Lakkaraju and Bastani, 2020; Goyal et al., 2023; Buçinca et al., 2021). Therefore, a key design objective is to steer clear of biasing the decision-making process and focus solely on encouraging justifications. This approach aims to assist novice reviewers in producing high-quality and original reviews without compromising the integrity of the evaluation process.